OpenStudy (anonymous):

What is the equation of the line that passes through (-2, 2) and (1, -4)?

OpenStudy (igreen):

Find the slope first. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-4-2}{1+2}\) Tell me what you get. @mr.cornist

OpenStudy (anonymous):

okay and sorry i dont kno how to use this

OpenStudy (igreen):

Subtract -4 - 2 and add 1 + 2, then divide.

OpenStudy (anonymous):

i got it thanks